OpenStudy (anonymous):

Change the equation into slope-intercept form. (y=mx+b) –2x + 3y = -6

OpenStudy (anonymous):

So for slope intercept, you're going to want to isolate y

OpenStudy (anonymous):

start by adding 2x to both sides to come up with 3y=2x-6

OpenStudy (anonymous):

Now divide both sides by 3 to get y alone, you'll come up with y=2/3x-2

OpenStudy (anonymous):

Oh sorry that was the last step lol y=2/3x-2 :)

OpenStudy (anonymous):

Okay, thank you for also showing me how to do it. I appreciate it ^-^

OpenStudy (anonymous):

You are very welcome, do you understand how to do it now?

OpenStudy (anonymous):

yes, and by y=2/3x-2 is that a fraction bar?

OpenStudy (anonymous):

Yes ma'am 2/3x is the same as saying 2 ____ x 3

OpenStudy (anonymous):

alright, thank again